OpenStudy (anonymous):

solve for a 4a+3c=-2a+5 answer in this format a=(__)/__

OpenStudy (anonymous):

4a+2a=5-3c 6a = 5-3c a = (5-3c)/6

OpenStudy (triciaal):

4a + 3c =-2a +5 6a = 5 - 3c a = (5 - 3c)/6

OpenStudy (anonymous):

4a+3c=-2a+5 4a+3c+2a=5 4a+2a=5-3c 6a = 5-3c a = (5-3c)/6

OpenStudy (anonymous):

@Shay17 your really helpful thanks :)

OpenStudy (anonymous):

yw : )